RRRG new website top logo resized

 

Restoration diary

Reports of restoration work on 50029 and 50030.

RRRG website overhaul

The old RRRG website was a cutting edge design... for the late 1990s.  As a result it has been sent to Doncaster Works for a total refurbishment and is expected back in traffic later this year.  Meantime we have set up this site to provide news, views and project updates until such time as the refurbishment work is completed.

Continue reading
3002 Hits
4 Comments

Working party 4-5 Feb 2012

We will be on site over the weekend of 4-5 February working on our locomotives.  Come along and say hello; we will be delighted to talk about our project to you.

Tags:
Continue reading
2537 Hits
0 Comments

July 24th 2011

Dave was on site all weekend as usual, and was attempting to loosen the bolts on the radiators with an air-tool, but discovered it seems to be broken, so he went to remove some of the brake rigging from the spare bogies, but found all the bolts seized. Wes Needle paid a visit with his wife Keren & their two children on Sunday to look in on progress and take the family for a ride on the train. As I was working on 50008 last Sunday & therefore missed the workparty, I went up today & fitted the second small board behind the field divert contactors that houses the ETH detection relay HDR, the ETH earth fault relay HEFR & it's associated resistor Z62. I wired Z62 & HDR, but while wiring HEFR, I seem to have mislaid a wire from one of the ETH interlock relays, so finding it is the first job next time I am there.

Tags:
Continue reading
3125 Hits
0 Comments

July 16th-17th 2011

It was quite a busy weekend with Dave, Chris, Tom, Steve & Sarah attending both days and Ian, Pete & Mark on site Sunday. Toby the dog, who could become un-official mascot with regular site attendance, was also there!! Ian & Pete continued with the quest of torqeuing up the big end bearing caps, only a couple left to do.

Tom repaired/replaced as required all the runner timbers below the droplights. This completes all the internal frame repairs in both cabs.  The potentiometer stand in Renown was straightened, welded, cleaned and refitted by Chris - it had been damaged and bent during the removal of the ETH/aux gens before RRRG bought the locos. Dave continued cleaning up under the no1 cab floor in Renown, also assisited others.

Steve & Sarah continued identifying cut cables in Renown concentating at the cubicle end. They opened up the Battery Isolation Switch to identify cables, the rear section of the switch was removed to assist this process.

Mark removed floor-mounted conduits from the generator area in Repulse. The conduits will be replaced with those removed from 50040 prior to its scrapping. The cleared floor was cleaned.

Tags:
Continue reading
2243 Hits
0 Comments

Missing posts gap

Unfortunately the posts in our restoration diary from July 2011 until February 2012 are missing, having been consumed by the Great RRRG Website Crash of 2011.  The posts that follow through to November 2012 are taken from our interim Wordpress site at http://renownrepulse.wordpress.com which was created to give us a web presence whilst we overhauled the main site www.renownrepulse.com which you are now reading.  Reporting activity will now continue on this site.

Continue reading
2559 Hits
0 Comments

June 26th 2011

Ian & Pete stretched another two pistons' worth of stretch bolts, before having to abandon this until a spanner can be bent, as the power unit internal structural webbing is now getting in the way of using the spanner properly. They then helped Dave loosen the bolts on the spare bogies to allow the traction motors to be removed sometime this year for overhaul. I test fitted the other two small boards that go behind the divert contactors, and found they fit easily. These two boards house the ETH overvolt relay HOVR & it's associated resistor Z24, the ETH voltage sensing resistor Z23, the ETH earth parallel resistor Z69, the ETH earth fault relay HEFR, & it's associated resistor Z62, and the ETH detector relay HDR. I bolted the left-hand board in place and wired in HOVR, Z23, Z24 & Z69.

Tags:
Continue reading
2853 Hits
0 Comments

June 20th-23rd 2011

Chris & Tom spent four days on site doing various jobs including repairing, straightening & painting the mounting pedestal for the load limiting pot, cabside corrosion repairs to No.2 cab of Repulse, collecting the spare auxiliary gen from Carnforth & taking it to a secure location, continuing to see if a new aux contact mechanism can be made for the reverser for Renown, creating shelving for me in the air filter area of Repulse and various other jobs.

Tags:
Continue reading
12479 Hits
0 Comments

June 12th 2011

Dave as seems to be the case most weekends at the moment, spent both days doing various jobs, then helped Ian & Pete stretch the stretch bolts on another three power unit big ends on a miserable Sunday when it rained continuously. I wired the auxiliary contacts on BMR1, GFR & OIR relays and the Z59 resistor on the board behind the fuse cabinet in Repulse.

Tags:
Continue reading
2979 Hits
0 Comments

June 4th-5th 2011

Dave was on site all weekend whilst Steve & Sarah spent Saturday afternoon and all day Sunday there. Dave carried out various work on brake rigging and some bodywork both on Repulse. He also looked into preparing the traction motors for lifting from the two spare bogie sets, I think the fixing bolts are proving very tough!!!


Steve & Sarah continued identifying the cut cables in Renown. The going was very slow this weekend as the cover to the water header tank level switch would not budge so in the end with help from Dave the whole level switch assembly was removed from the tank. Once the whole assembly was removed the cover was prised away to reveal the cable connections. The two level switch cables were identified and labelled. The level switch operation was tested and found to be good and therefore the whole assembly was reinstalled.


A further 2 cables was identified at the Fire Alarm Relay No1. All the cables were retested at the AWS Relay Box to prove we had identified these.
Steve and Sarah took an hour out on Sunday to meet up with John Patrick - Peak Rail Safety Officer to complete their Personal Track Safety certification.

Tags:
Continue reading
2861 Hits
0 Comments

May 21st-25th 2011

Dave, Chris & Tom spent a marathon five days on site. Dave sorted the spare brake rigging and began cleaning & un-sticking various parts, Tom continued working on the interior wood surrounds on the quarter-lights & drop-lights in 30's No.2 cab & Chris fitted the three remaining sections of 185mm cables to the field divert contactors, as he & I had decided a fresh pair of eyes might be able to devise a solution from pictures he had taken of the cables in 50050. He has also taken the auxiliary contacts off the 'proper' 50 reverser, to see if he can fabricate a new mechanism for the second reverser for Renown & begun cleaning debris from the No.1 end clean air compartment.

Tags:
Continue reading
2949 Hits
0 Comments

May 15th 2011

Dave spent the weekend on site clearing up the stores shed, then on Sunday helped Ian & Pete stretch 12 more stretch bolts on three big ends on the power unit, having made a new tool during the week. Having forgotten to reconnect the rad fan fuse last week (despite posting that I had completed the wheelslip board), I did so, then fitted the small relay board behind the fuse & mcb cabinet. This board has the Overload Interlock relay OIR, Generator Field relay GFR, No.1 motor blower relay BMR1 and the lighting time switch load resistor Z59. Two of the mounting bolts again required loosening slightly to fit the board in the somewhat tight space, and the resistor was removed to ease fitment. The lugs on the large cables to the GFR & BMR1 were cleaned and they were bolted on to their connections on the rear of the board, which again was time consuming due to the confined space to wield spanners.

Tags:
Continue reading
2476 Hits
0 Comments

May 1st 2011

Dave spent the weekend on site having a general tidy-up of the coach and container, then helped Ian & Pete torquing the big-end stretch bolts until the tool broke! After I had returned the wheelslip relay board to my house after the previous workparty, I removed the contact fingers from the relays and cleaned them (not the easiest of tasks as the fixing screws are tiny & probably hadn't been removed since the relays were originally made), straightened two of the contact fingers on one relay. They were re-assembled and fitted back onto the paxolin board with new brass cable connection bolts to replace to broken originals, checked for operation by dabbing a 110v DC supply onto the coil contacts and were found to be still working. The associated three 10k ohm resistors were also removed, cleaned, checked with a meter to see they were still working, and refitted. The brass rad fan fuse mounting pillars were missing, so they were taken from the spare broken board, cleaned & fitted. The mounting pillars for the plastic relay covers were removed as they were rusty, & were de-rusted, painted & refitted. The filthy relay covers were cleaned and refitted. Today, the completed board was mounted in the elec cubicle in Repulse, but required a small piece cutting off one corner as some cables that go past the board were hard against one edge, and two of the mounting bolts needed to be loosened as they didn't quite line up with the mounting holes. The board was then fitted and bolted in, the rad fan fuse was fitted & the external connections have been made.

Tags:
Continue reading
3278 Hits
0 Comments

March 20th 2011

Chris & Tom Thorn & Dave again spent the weekend on site starting to remove the removable rad grille section and the air filter frame in the passageway, to get at the motor blower inside. Chris B joined them today & by the end of the day, they had unbolted & removed the blower snail housing out through the hatch for cleaning. It is hoped the motor itself will follow soon, but due to it's weight, it will probably have to be lifted out through the roof. Ian & Pete removed all the cam followers from the power unit's B side for cleaning & to blow out the oilways with the compressor. I rewired part of the rear of the cleaned wheelslip relay board, using new brass nuts & bolts, as the original cable mounting bolts had all been sheared off sometime ago. Once the wheelslip resistors have been cleaned, and I have verified the relays are still working, the board will be reassembled & can then be fitted into the elec cubicle.

Continue reading
3546 Hits
0 Comments

March 6th 2011

Chris T continued sorting & tidying in both coaches, and applied the 'Smartwater' security stickers to the coaches & both locos, Tom continued work in the no.1 cab of Repulse, Ian & Pete continued fitting the water jacket flexible elbows from the engine block to the cylinder heads, but found two that seemed reluctant to fit, so left them for next time. I tested all the wheelslip resistors with my meter to make sure they were still intact, then removed two, without breaking the mounting bolts, to go with the one I had previously removed, and brought them home to fit to the board I took home last time once cleaned. Chris T thinks he has found a supplier of the Londex wheelslip relays, in case the original ones prove unusable & I will investigate this further. I persuaded the recalcitrant mounting bolts I mentioned last time to align with another relay board's mounting holes with elbow grease and a nylon hammer, test fitted the board, dismantled and cleaned the auxiliary contacts of the relays on said board and started cleaning the somewhat soiled 35mm cable lugs that connect to them.

Tags:
Continue reading
2847 Hits
0 Comments

February 20th 2011

Dave spend the day cleaning the brake chain & mechanism in the no 1 cab of Renown, Ian, Pete & Chris B removed the 12 remaining fuel pumps to be sent for overhaul & I tidied some more completed cable runs in the cubicle of Repulse, started to bend the fixing bolts for the board mentioned last time & examined the three boards with the wheelslip relays & their resistors. There are three wheelslip relays & three resistors on each board & one board is required per loco. One board seems to have been removed from a 50 rather roughly, as the one corner is broken off where the one fixing hole was, so that one is not usable. The other two boards are intact but someone had tried to remove the resistors before we acquired the boards, and sheared off their fixing bolts level with the boards in the process. Many of the cable fixing bolts for the relays are similarly sheared and can be replaced with brass ones, but the resistor bolts will need further investigation, so I brought one board home to see if I can devise a way to re-secure them.

Tags:
Continue reading
3239 Hits
0 Comments

January 30th 2011

Dave, Ian & Pete torqued the remaining 10 cylinder heads down to 600 ft lbs, a somewhat lengthy task, having to remove and replace two of the large head fixing studs which had damaged threads. Dave had spent the weekend on site again & during Saturday had fitted new flooring in the no2 cab of Repulse. Chris B unbolted three fuel pumps. Tim & Dom had a good tidy up around our locos and coach, and removed old cab flooring round to the bonfire site, and I fitted and wired the battery charge diode, tidied up some sections of cabling with cable ties & Chris B helped me thread a long section of 16mm cable into the cubicle which connects one of the circuit breakers to terminal bar 7. With that last section of cabling fitted, all the fuses and mcb's are fully wired, and the small relay board that goes in front of them can now be fitted.

Continue reading
3134 Hits
0 Comments

January 16th 2011

Tim & Dom did housekeeping duties around our vehicles, then started hoovering accumulated debris from under the cab floors of Renown & the rad compartment passageways. Ian & Pete fitted the last three big end bearing caps into the power unit. Dave started to free off the siezed handbrake mechanism in the no1 cab of Renown and cleaned & painted more cab parts. Now Steve & Sarah have recovered from their car crash enough to be more active, they visited site & I discussed various options with Steve for replacing severed cables inside Renown. They then removed some of the flooring alongside the elec cubicle and by the brake frame & started removing accumulated debris from under the floors. We had a quick look to see if any of the severed cables could be matched with their 'other' ends, and we think we may have identified those to the DSD relays. I test fitted a small relay board inside the electrical cubicle of Repulse, but the fixing bolts need bending a little to bring them back into alignment with the board's mounting holes. I also fitted the Battery Charge Diode. Kev Hand made a surprise visit & returned both reversers he was storing for us, as we are now in a position to permanently fit the proper one into Repulse.

Tags:
Continue reading
3671 Hits
0 Comments

December 11th 2010

Dave Rolfe continued measuring, cutting and fitting flooring into the no.1 cab of Renown. Chris Bodell painted some more cab fittings & panels, and varnished the completed flooring in no.2 cab, Ian & Pete fitted two big end bearing caps now the suppliers have sent Ian the nuts for the stretch bolts that they were out of stock of. They were in the process of fitting a third cap but one of the four stretch bolts seemed unwilling to fit, so they had to remove the cap to see what the problem was. Nothing was obviously amiss, but by then the light was fading, so they didn't refit the third cap. My two GWR mates (Matt & Roger) who had visited a few weeks ago again came up, this time armed with some 185mm through-crimps and a large hydraulic crimper, so in the morning we did a test joint to see how it looked, and the result was encouraging. After crimping both sides of the crimp, a layer of heat-shrink sleeving was applied and heat-shrunk. A large amount of insulation tape was then wrapped round the crimp over the sleeving, then a second layer of heat-shrink sleeving was applied and shrunk. We used the group's 1000v Megger to check if there was any leakage to earth, and the Megger result was a full-scale needle deflection to the Megger's capacity of 2 Mega Ohms (2 million ohms). Any reading over 1 megaohm is considered good, and as the Megger only registers up to 2 Meg, the actual reading would therefore have been higher, which means this type of crimp, which is rated at 11,000 volts, should be more than adequate for both the severed main traction cabling and the main ETH cabling, as the maximum expected traction voltage is around 1000 volts. If I can find someone with a higher-capacity Megger, I will repeat the insulation test. Joining the cables in this way will remove a significant amount of time, effort and expense from the work needed to complete Repulse. After lunch, we crimped the last couple of 16mm cables, connected them, then Matt connected four cables to resistors on top of the cubicle which I am too much of a shortarse to reach.

Tags:
Continue reading
5571 Hits
0 Comments

November 14th 2010

Dave started fitting new flooring in the no.2 cab of Renown, Tom continued work in the no 1 cab of Repulse, Chris T & Tim continued loading items into BG2 and rebuilding the horn air feeds in 30 with parts taken from the cab of 50037, Chris B spent the day painting power unit coolant and oil pipes, Ian & Pete refitted the last cylinder head and began fitting some of the water piping, Dom lifted the remains of the rotten flooring in the no.1 cab of Renown & cleaned out the accumulated debris underneath & Mark brought some recently acquired spares to site & undid the remains of one set of the severed generator cable conduits in the clean air compartment. Two ex-GWR colleagues of mine visited today, partly for a look at progress on 30, partly to advise me on possible ways to join the severed motor cables under the cubicle floor without having to rewire with expensive new cables, and partly to lend me their hydraulic crimper to put new lugs on a few more 16mm cables.

Continue reading
2736 Hits
0 Comments

October 31st 2010

Ian & Pete fitted three pistons and two heads during the day, Chris B continued working in the no.2 cab of Repulse refitting internal panelling, and I fitted several 16mm & 35mm cables I had crimped new lugs on, where the originals had been cut off & fortunately they were still long enough to re-fit them without incident.

Tags:
Continue reading
2700 Hits
0 Comments